OpenStudy (anonymous):

how would i graph this? y=-1/2x-3

OpenStudy (anonymous):

a few ways. but an easy one would be this : it is linear, so you will definitely get a line. pick any two points on the line and connect them. pick easy ones @ x = 0, y = (-1/2)(0) - 3 = -3 so (0, -3) @ x = -6 , y = (-1/2)(-6) - 3 = 3 - 3 = 0 so (-6, 0) just plot those two points and connect them!

OpenStudy (anonymous):

thanks :)